What are the odds of rolling double six in dice-throwing?
This can be considered two independent Bernoulli events, so the probability of the whole is the product of the individual events, or (1/6)(1/6) = 1/36.
What is difference between uncertainty and probability?
Uncertainty is the uncertainness of any happening which may occur but probability is the ratio of which the happening will occur.The number of happening can also be included in the probability section but not in uncertainty.Both the words are roughly opposites of each other.
What is the formula for probability fraction?
Usually, a fraction is already a probability. Like if you roll a die, the probability that you will get a 1 is 1/6. So your probability would be your fraction.
What is a subset of sample space?
A subset of sample space is taking a sample from that sample space.
What is the probability of drawing a 6 and a spade?
Probability of drawing a spade is 13/52.
Probability of also drawing a 6 is 3/39. Therefore it would be 13/52 x 3/39= .0192
What is the probability of obtaining exactly 4 heads in 6 flips of a fair coin?
The probability of a head in one flip is 1/2.
The probability of HHHHTT is (1/2)6 = 1/64
The possible correct flips are HHHHTT, HHHTHT, HHTHHT, HTHHHT, THHHHT, HHHTTH, HHTHTH, HTHHTH, THHHTH, HHTTHH, HTHTHH, THHTHH, HTTHHH, THTHHH, TTHHHH, each with a probability of 1/64.
Total probability is 15/64.
Is the area under the standard normal distribution to the left of z equals 0 negative?
The Z value is negative, but area is always positive.
What is the probability of rolling double sixes with 2 rolls of the die?
Rollin for dubs- check em
1/6 chance for rolling the first six. 1/6 chance for rolling the second one (independantly). when doing compound probability problems like this you multiply the chances of each soooo
1/6 x 1/6 = 1/36 chance of rolling dubs sixes (or any other number for that matter)
hell no nigaa
How many possible outcomes of tossing a coin 10 times have at least 2 heads?
There are 210 total possible outcomes from flipping a coin 10 times.
There is one possible outcome where there are 0 heads.
There are 10 possible outcomes where there is 1 head.
So there are 210 - 11 possible outcomes with at least 2 heads.
(1013)
What is the probability of answering a 9 question true false test correctly if you randomly guess?
The probability of getting the first answer correct is 1/2
The probability of getting the first two correct is 1/2 * 1/2 = 1/(22)
The probability of getting all 9 correct is 1/(29) = 1/512 which is just under 0.2%
What is the probability of choosing a marble that is not blue in problem 2?
if there is a jar containing 5 red marbles 6green and 4 blue what is the probability off not chossing a blue marble
What does at least as large as mean?
It means the size is the same or larger than the item being referred to.
What probability value would be needed to complete the following probability distribution?
You should be given p(x) values such as 0.09 0.19 0.14 0.29 you just add these values to get 0.71 subtract 0.71 from 1 to get 0.29 is the answer
What is the probability of getting a white offspring from this cross?
We would need to know the situation to respond to this question.
What is an example of Empirical Probability?
Empirical means by observation, so empirical probability, or experimental probability, is the probability that is observed in a set of trials. For example, if you flip a coin ten times and get seven heads, your empirical probability is 7 in 10. This is different than the theoretical probability, which for a fair coin is 5 in 10, but that result will only be approximated by the empirical results, and then only with a larger number of trials.
What does it mean if the standard deviation is equal to the mean?
A large standard deviation indicates that the distribution is heavily weighted far from the mean. Take the following example:
{1,1,1,1,1,19,19,19,19,19} Mean is 10 and StDev = 9.49
Now look at this data set: {5, 6, 7, 8, 9, 11, 12, 13, 14, 15}
Mean is still 10, but StDev = 3.5
How much steel required for square meter in roof slab?
for residential building 1 sqm steel required =8kgs
1 in 32. Each correct answer has to rely on the previous answers all being correct. By random chance getting five heads in a row or picking black or red correctly, on a wheel with no green slots, five times in a row is 1 chance in 32 attempts.
27/35
The binomial and Poisson distributions are examples of discrete probability distributions?
discrete & continuous
What is the probability of chossing 12 marbles from a jar in numercial order?
The probability of picking the #1 marble on the first draw is 1/12.
If you've done that, then the probability of picking the #2 marble on the second draw is 1/11.
If you've done that, then, the probability of picking the #3 marble on the third draw is 1/10.
If you've done that, then, the probability of picking the #4 marble on the fourth draw is 1/11.
etc.
etc.
So the probability of doing all of them in sequence is (1/12) x (1/11) x . . . x (1/1).
That's exactly the reciprocal of (12!). According to my $1.49 calculator, your chances
of success without peeking amount to about
0.00000020877 percent (rounded)
Not a smart bet.